OpenStudy (anonymous):

What is 1 over 7 of 42 ?

OpenStudy (anonymous):

pls write this .;)

OpenStudy (math&ing001):

\[\frac{ 1 }{ 7 }49\]

OpenStudy (math&ing001):

42*

OpenStudy (anonymous):

Look : we have an equation : 1/7 x = 42 Right ???? @mario_4567

OpenStudy (charlotte123):

it would be 6 http://www.wolframalpha.com/input/?i=1%2F7+of+42

OpenStudy (math&ing001):

@E.ali actually it's \[\frac{ 1 }{ 7 }*42=\frac{ 42 }{ 7 }=6\] He didn't mention any equation

OpenStudy (charlotte123):

hurray!

OpenStudy (math&ing001):

:p

OpenStudy (anonymous):

ahahahahahhahahah !!!! ok ok oko ko ko ko ko ko ko ko koookokkkokok !!!!

OpenStudy (charlotte123):

LAWL